OpenStudy (anonymous):

A tumor is a collection of cancerous cells. Suppose each cancerous cell has a radius of 5 times 10^{-3} cm, and that the doubling time for a population of cancerous cells is 50 hours. If the tumor has an initial volume of 0.1 cm^3 (For simplicity, assume that the tumor is entirely made of cells, that there is no space between them.) Round your answer up to the next whole number. Find a formula for V(t), the total volume of the cancerous cells, after t hours. V(t)= How many hours will it take before the volume of the tumor reaches 1 cm^3?

OpenStudy (anonymous):

you are given your initial volume of 0.1 cm^3, so that is your 'a' you are given that at t = 20 hours the volume doubles to 0.2 cm^3. use this and the initial volume to find b. Then put it into a*b^t form. don't round, use the exact answer you found b to be. Meaning when you solve for it using algebra, don't plug anything into your calculator. As soon as you use your calculator you are rounding.

OpenStudy (anonymous):

This is the reply from my instructor but I don't follow

OpenStudy (anonymous):

if total volume/ volume of a single cell = # of cells. Then total volume = (# of cells)* volume of a single cell now the # of cells is an exponential equation and the volume of a single cell never changes. So, you will need to use your initial # of cells to find your 'a' and then the point of doubling to find your 'b' and then you have total volume = (a*b^t)* volume of a single cell.

OpenStudy (whpalmer4):

Can you write the equation for the number of tumor cells as a function of \(t\)?

OpenStudy (whpalmer4):

note well that we've gotten two different doubling times specified here, 20 hours and 50 hours!

OpenStudy (anonymous):

This is as far as I got on it. 4/3*pi*r^3=5.23598*10^-7 1.91*10^5=191,000

OpenStudy (anonymous):

If the tumor has an initial volume of 0.1 cm^3, how many cancerous cells does it contain? (For simplicity, assume that the tumor is entirely made of cells, that there is no space between them.) Round your answer up to the next whole number. The answer above is for this first part of the problem

OpenStudy (anonymous):

I am kind of lost on how to get V(t) given the initial size and find b with a

OpenStudy (whpalmer4):

190,986 cells is what I get, but what's 14 malignant cells between friends? :-)

OpenStudy (whpalmer4):

So let's call that \(P_0\) for the initial population of tumor cells. Now we want to develop an equation that will give us \(P(t)\) or the population of tumor cells at time \(t\).

OpenStudy (whpalmer4):

did you decide if you're going to use 20 or 50 hours as the doubling time?

OpenStudy (anonymous):

20

OpenStudy (whpalmer4):

okay, so \(t_{dbl} = 20\) We can write our population equation as \[P(t) = P_0 (2)^{t/t_{dbl}}\] At \(t=0\), it reduces to \[P(0) = P_0(2)^{0} = P_0\] At \(t = t_{dbl}\), it becomes \[P(t_{dbl}) = P_0(2)^{t_{dbl}/t_{dbl}} = P_0(2)^1 = 2P_0\] Do you agree those both square with your understanding of the scenario

OpenStudy (anonymous):

Is double basically 2 and I would replace dbl with 2?

OpenStudy (anonymous):

Well, replace it with 20 as our doubling time

OpenStudy (whpalmer4):

\(t_{dbl}\) is just the time it takes for the population to double

OpenStudy (whpalmer4):

you replace it with 20 hours if you're going with 20 hours as the doubling time

OpenStudy (anonymous):

Ok, so we have the population equation and I have substituted 20 into the original equation for dbl

OpenStudy (whpalmer4):

uh, what is the "original equation"?

OpenStudy (anonymous):

Oh, into our population equation

OpenStudy (anonymous):

Sorry, there is no original

OpenStudy (whpalmer4):

so, the volume equation is just the population equation times the volume of a single cell, right?

OpenStudy (anonymous):

yes, that is right

OpenStudy (anonymous):

I have to come up with the formula to satisfy Find a formula for V(t), the total volume of the cancerous cells, after t hours. V(t)=

OpenStudy (whpalmer4):

Okay, so write out that equation and evaluate it at \(t = 0\). Then set \(V(t) = 2V(0)\) and solve for \(t\)

OpenStudy (anonymous):

That is where I am having the problem... don't know what equation to write out and what values to use

OpenStudy (whpalmer4):

Whoops, reread the problem, never mind that last post! You want to set \(V(t) = 1 \text{cm}^3\) and solve for \(t\)

OpenStudy (whpalmer4):

Look, didn't we just agree that the volume equation is the size of a single cell * P(t)? What more do you need?

OpenStudy (anonymous):

Yes, so in order to find t, I would set V(20) = 1cm^3^t/20 (confused at this part)

OpenStudy (whpalmer4):

the only place you get to use 20 is where I have written \(t_{dbl}\). It is a constant in the equation.

OpenStudy (whpalmer4):

please, write out the volume equation as you understand it for me.

OpenStudy (anonymous):

ok, V=Po(2)^t/20

OpenStudy (whpalmer4):

you need some extra parentheses in there to make it say what you mean if you aren't going to format it... V(t) = P0 (2)^(t/20) so, V(t) at the moment of interest is 1 cm^3 1 cm^3 = P0 (2)^(t/20) \[1 = P_0 (2)^{t/20}\]\(P_0\) is a constant you can put in once you've rearranged to solve for \(t\); do you know how to solve that equation for \(t\

OpenStudy (whpalmer4):

equation for \(t\)?

OpenStudy (anonymous):

no, I got part of the volume equation right, but not sure how to solve for t

OpenStudy (whpalmer4):

don't remember your properties of logarithms, eh? \[\frac{1}{P_0} = 2^{t/20}\] Take the log of both sides: \[\log\frac{1}{P_0} = \frac{t}{20}\log 2\]because \(\log u^a = a\log u\) \[\frac{\log \frac{1}{P_0}}{{\log 2}} = \frac{t}{20}\]

OpenStudy (anonymous):

it's funny that you should say that... I failed my last exam on logarithms (it's one of the areas I am not so good at)

OpenStudy (whpalmer4):

Actually, we're going from a volume of 0.1 cm^3 to 1.0 cm^3 which is a factor of 10. Because the volume varies directly with the population, we could just find the time for the population to grow by a factor of 10. \[P(t) = 10 P_0 = P_0 2^{t/20}\]\[10 = 2^{t/20}\]\[\log 10 = \frac{t}{20}\log 2\]\[\frac{1}{0.30103}*20 = t\]

OpenStudy (anonymous):

thanks for staying up late to help me with this problem... sadly I have been working on it off and on for a couple hours

OpenStudy (whpalmer4):

my equation above I forgot to multiply by the volume of a single cell! the second works much better :-)

OpenStudy (anonymous):

so t = 66.438561

OpenStudy (whpalmer4):

should have been \[1 = V_{cell} P_0 (2)^{t/20}\]where \(V_{cell}\) is the value you worked out for a single cell's volume Yes, that's the correct value for \(t\)

OpenStudy (anonymous):

or just 66.44

OpenStudy (anonymous):

I never know if I should round up to 2 or 4 decimal places

OpenStudy (whpalmer4):

Here's a plot, showing each doubling (1st 3 vertical lines) and then the final arrival at V=1 cm^3 at t = 66.44 hours

OpenStudy (anonymous):

you are probably in a higher level math such as calculus or beyond

OpenStudy (whpalmer4):

uh, well, I was the last time I took a math class, but that was many years ago :-)

OpenStudy (anonymous):

I have to take precalc for my b.s. in software development

OpenStudy (whpalmer4):

well, you'd probably get more use out of a discrete math class than precalc, but depending on what you'll be developing, no lasting harm should come from exposure to this material :-)

OpenStudy (anonymous):

so now that I got t, I plug that in to get the formula for V(t)

OpenStudy (anonymous):

I am still on that part of the question :)

OpenStudy (whpalmer4):

no, you already have the formula for V(t). It is the volume of a single cell * P(t)

OpenStudy (whpalmer4):

if you want a nice general formula, \[V(t) = \frac{4}{3}\pi r^3 P_0 (2)^{t/t_{dbl}}\]where \(r\) is radius of a single cell, \(P_0\) is the initial population of cells, \(t_{dbl}\) is the doubling time

OpenStudy (anonymous):

that is going back to this formula: \[V(t)=2^(t/20)\ ]

OpenStudy (anonymous):

oh, I was sort of on the right track

OpenStudy (anonymous):

this problem is difficult because I am trying to wrap my head around volume of population of cancer cells... an abstract idea

OpenStudy (whpalmer4):

Actually, it could just as easily be the weight of the tumor — the only change would be that we'd find the weight of an individual cell and multiply by that instead of the volume of an individual cell.

OpenStudy (anonymous):

yes, it could be anything... so given the general formula, it is still does not accept that answer although I am not sure if I put in the equation correctly

OpenStudy (anonymous):

Find a formula for V(t), the total volume of the cancerous cells, after t hours. V(t)=

OpenStudy (anonymous):

I put in 4/3*pi*r^3Po(2)^t/20, maybe I need to simplify this when I put in my answer

OpenStudy (whpalmer4):

you need to substitute your values in... you're going to have \[V(t) = <number>2^{t/20}\] another possibility is that you are supposed to be using 50 hours as the doubling time (remember, that's what it says in the initial problem statement)

OpenStudy (anonymous):

oh, so that is what the instructor is talking about where I was given the values for initial value and t=20

OpenStudy (whpalmer4):

yes

OpenStudy (anonymous):

ok, 4/3*pi*r^3 is just the formula for volume in general

OpenStudy (whpalmer4):

yes

OpenStudy (anonymous):

but it doesn't look like I used that formula in the problem at all

OpenStudy (anonymous):

except to find the first value

OpenStudy (whpalmer4):

you used it to find the volume of a single tumor cell.

OpenStudy (whpalmer4):

the volume of a single tumor cell is part of the equation

OpenStudy (whpalmer4):

it is also a constant for this problem

OpenStudy (anonymous):

ok so probably still missing something but for the formula I have V(t)=0.1^(t/20)

OpenStudy (whpalmer4):

yes, you're missing the 2!

OpenStudy (whpalmer4):

and the initial population

OpenStudy (anonymous):

so the formula we are working with is Ae^kt

OpenStudy (anonymous):

I remember some of this

OpenStudy (anonymous):

ok, second part is correct V(t)=0.1(2)^t/20

OpenStudy (anonymous):

that is what they asked for in V(t)=

OpenStudy (anonymous):

one last part to all of this How many hours will it take before the volume of the tumor reaches 1 cm^3? Round your answer up to the next whole number.

OpenStudy (anonymous):

so hours = t

OpenStudy (whpalmer4):

We solved that already! Remember the stuff with the logarithms?

OpenStudy (whpalmer4):

t is in units of hours, yes

OpenStudy (anonymous):

log10 = t/20 log2

OpenStudy (anonymous):

1/0.30103 * 20 = t

OpenStudy (whpalmer4):

yes, go on

OpenStudy (anonymous):

t = 66.44

OpenStudy (whpalmer4):

close enough for government work...

OpenStudy (anonymous):

so we already solved that at the same time as finding V(t)=

OpenStudy (anonymous):

that makes things easier

OpenStudy (anonymous):

66.44 is not going through

OpenStudy (anonymous):

Round to the nearest whole number, t=66.438561

OpenStudy (anonymous):

67 = t

OpenStudy (whpalmer4):

no!

OpenStudy (whpalmer4):

nearest whole number is not 67

OpenStudy (whpalmer4):

66.438561-67 = -0.561439 66.438561-66 = 0.438561 which number is closer, 66, or 67?

OpenStudy (anonymous):

66

OpenStudy (anonymous):

ok, I put in 66 and still not what they are looking for... going back to the value of t

OpenStudy (whpalmer4):

that answer is going to depend on what you used for the doubling time, of course.